simplify log[7] 7^5
Answer by jessica43(140)

To solve this problem, you are going to need to use the base b logarithm formula to rewrite the problem:
log[base b]a = c means the same as b^c = a
So use this in your problem:
log[7] 7^5 = x
a = 7^5, b = 7, c = x
Now rewrite it as an exponent:
b^c = a
7^x = 7^5
x = 5
So your answer is x = 5.